OpenStudy (anonymous):

a positive interger is 34 more than 30 times another. Their product is 22788. Find the two integers.

OpenStudy (anonymous):

Let x = one of the integers and y = the other. x-34=30y x+y = 22788 Solve the equations.

OpenStudy (anonymous):

Yea I thought that too but I was wrong. Here is how to solve. X= one interger 30x+34= the other positive interger X(30x+34)=22788 Solve with Quadratic equations.

OpenStudy (anonymous):

that would work as well. I just used the equations I noted and got x=22788-y and then substituted that into the first one